OpenStudy (anonymous):

The distance between the store and jade's house is 500 meters. She walked 1m/s for 5 minutes , How many meters are left? The distance between the store and jade's house is 500 meters. She walked 1m/s for 5 minutes , How many meters are left? @Mathematics

OpenStudy (saifoo.khan):

5 min = 5*60 = 300 secs. speed = distance/ time 1 * 300 = distance distance= 300m.

OpenStudy (saifoo.khan):

So, 500 - 300 = 200m left.

OpenStudy (anonymous):

Thank U , I Had Forgot The Equation .